OpenStudy (anonymous):

Given h(x) = 3x/-x-1 find h^-1(x)

OpenStudy (anonymous):

ok that was wrong solve \[x=\frac{3y}{-y-1}\] for \(y\)

OpenStudy (anonymous):

Multiply by -y-1?

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{h(x) = \dfrac{3x}{-x-1}\\~\\ y=\dfrac{3x}{-x-1}\\~\\ -y=\dfrac{3x}{x+1}\\~\\ -yx-y=3x\\~\\ -y=3x+yx\\~\\ -y=x(3+y)\\~\\ x=\dfrac{-y}{y+3}\\~\\ \text{swap x and y}\\~\\ y=\dfrac{-x}{x+3}\\~\\ \color{red}{h^{-1}(x)=\dfrac{-x}{x+3}}\\~\\ }\end{align}\)